# Daily Land Surface Temperature Reconstruction in Landsat Cross-Track Areas Using Deep Ensemble Learning With Uncertainty Quantification

### v1

@@ III-A2 London
To demonstrate the generalization of the proposed framework, London is selected as the second study case. London is the largest city in the U.K. and Western Europe, with 8.8 million residents in 1,572 km2 area (5,690 per km2). Apart from a small portion in the southeast, the majority of London falls within the Landsat cross-track areas (Path 202 Row 024, Path 201 Row 024). London has two levels of small geographical units (with 983 Middle Super Output Areas and 4,835 Lower Super Output Areas) that are finer than 1 km2 – the MODIS resolution (Figure 5). These geographical units have mortality and other socioeconomic and health statistics that can be used for social good, but as a result of the lack of high resolution temperature data, existing studies had to settle on using coarse resolution temperature data that limits their capability [22, 36, 37].
III-A3 Hong Kong
@@
Hong Kong is selected as the third study area to demonstrate the generalization of the proposed framework and method across different continents and climate zones. Hong Kong is a coastal city in Asia, with a land area of 1,073 km2, housing 7.5 million population (6,989 per km2). Similar to New York City and London, most urban areas are also situated within the Landsat cross-track areas (Figure 5). The city maintains a large portion (75%) of undeveloped lands, further intensifying population density. As a high-density and subtropical city, Hong Kong faces various extreme climate events, in need of high spatial resolution temperature data.
@@
Figure 5: Meteorological stations of the three cities, overlapped with annual mean temperature and geographical unit boundaries (NYC: 135 merged zip codes, London: 983 Middle Super Output Area, Hong Kong: 452 District Council Ordinary Election Constituency Boundaries). New York City: 7 stations. London: 6 stations. Hong Kong: 23 stations.
III-B Data and Preprocessing
III-B1 Landsat Land Surface Temperature
@@
We obtained all 2023 Landsat 8 and 9 Collection 2 Level-2 science products and used the analysis-ready LST data from the thermal band (Band 10). NYC is completely located within two Landsat scenes (Path 013 Row 032, and Path 014 Row 032, Figure 2), with a total number of 92 scenes. London is completely covered by one scene (Path 201 Row 024), and has its majority covered by another (Path 202 Row 024), with a total number of 86 scenes. Hong Kong has most of its urban areas covered by two scenes (Path 121 Row 045, Path 122 Row 044), with a total number of 91 scenes.
@@
Before conducting the experiments, we went through a series of preprocessing, mainly on cloud masking and removal. We first checked the cloud band from the Level-2 products and masked out all data labeled as unclear. The cloud masks are generated from the Fmask algorithm and are not 100% perfect as noted from data description. We then double-checked visually with the imagery and cloud masks using a 3$\times$3 split. Within each split, if the pixels were mostly flagged as clear but were in fact covered by clouds, we masked out all the pixels within the split. In some cases, within each split, all pixels within the split appeared clear and most pixels were flagged as clear, but some pixels were flagged as not clear (mostly at coastline), we flagged this split as clear. After preprocessing, we have three types of data: dates under clear-sky situations, dates under partially-cloudy situations, and dates without valid observations (100% cloudy or no Landsat overpass).
III-B2 ERA5 Reanalysis LST Data
@@
III-B3 Annual Mean Spectral Reflectance
In order to link valid-observed pixels to cloud-covered pixels, the proposed method DELAG uses Gaussian processes to model the residuals between observed LST and ATC-reconstructed LST. The residuals are related to ground objects, which can be distinguished through spectral reflectance. We obtained the mean surface reflectance from 2022-2024 for four spectral bands (red, green, blue, and near infrared) from the Sentinel-2 products. The data is then resampled and matched with the Landsat ones. We obtained through Sentinel-2 instead of Landsat because Sentinel-2 has a higher revisit frequency, but the influence should be negligible. We also added the x and y coordinates, along with the spectral bands, as the features used in Gaussian processes.
III-B4 Near-surface Air Temperature Data
There is no in-situ LST site within the three cities, thus a direct comparison with in-situ LST in our study areas is not possible. However, given that most real-world applications require near-surface air temperature instead of LST, and many seamless LST data will be used for generating near-surface air temperature, we can indirectly validate the reconstruction with near-surface air temperature data (details in method Section IV-B3). This is an approach commonly used in the existing literature [18, 31]. For this purpose, we obtained air temperature data from meteorological stations. For New York City, we obtained data from 7 stations from the Automated Surface/Weather Observing Systems (ASOS/AWOS) maintained by the National Weather Service. For London, there are 6 meteorological stations. These data were obtained through Meteostat, an open source Python library that archives meteorological data from public sources. For Hong Kong, we obtained air temperature data from 23 stations from the Hong Kong Observatory. The locations of the meteorological stations in the three cities can be found in Figure 5.
IV Methodology
IV-A DELAG: Deep Ensemble Learning With Annual Temperature Cycles and Gaussian Processes
@@
Overall the proposed method DELAG can be expressed as
@@
IV-A1 Enhanced Annual Temperature Cycles
@@
IV-A2 Gaussian Processes to Model LST Residuals
@@
[…]
@@
IV-A3 LST Residuals for Cloud-Covered Pixels
@@
IV-A4 Deep Ensemble Learning to Find Model Solutions and Achieve Uncertainty Quantification
Given the complexity of the model as a non-convex function, it is very difficult, if not impossible, to find the global optimized set of parameters. In the PyTorch deep learning framework, we use the Adam optimizer (a variant of stochastic gradient descent) [40] to minimize the L1 loss to find the optimized solutions in the first step. In training the model, the training loss will eventually flatten and fluctuate, but cannot reach a global minimum. In Figure 6, near the end of training, 334 out of 425 snapshots all fall below the smallest contour level (loss$<$0.06). At this phase of the training, all snapshots are acceptable optimized solutions of the model. With each iteration, the parameters update again, resulting another snapshot of model parameters. The disagreement drawn from these snapshots of parameters is large enough, and utilizing the disagreements among these snapshot solutions has become a popular method to achieve ensemble learning or quantify uncertainty in deep learning [41, 42, 43]. Ensemble learning to achieve more robust and reliable prediction is also common in weather predictions and remote sensing [44, 45].
Figure 6: Demo showing the loss surface and the model’s trajectory in training. As training progresses, the loss value becomes sufficiently small, with 334 out of 425 snapshots falling below the smallest contour level of 0.06. With slight variations in these solutions, leveraging all snapshots as an ensemble allows the proposed DELAG method to deliver robust predictions and provide uncertainty quantification.
@@ IV-B Validation Strategies
IV-B1 Validation With Landsat Clear-Sky Observations Under Real-World Cloud Patterns
@@
IV-B2 Validation With Partially-Cloudy Data
@@
IV-B3 Validation Via Estimating Near-Surface Air Temperature
@@ IV-C Training Setup
The experiments were conducted on Python using the PyTorch framework [49]. For ATC fitting, we used the Adam optimizer with a learning rate of 0.1 for 1200 epochs. The 200 ensemble predictions were obtained through every 4 epochs in the final 800 epochs. For Gaussian processes, we used GPyTorch [39], a popular GP framework built upon PyTorch. We used a mini batch of 1024 and the variational inference with 512 inducing points for fast GP computation. The training was conducted with the Adam optimizer with a learning rate of 0.05 for 50 epochs and then of 0.005 for 10 epochs.
V Results and Analysis
V-A Experimental Setup
@@
We conducted three sets of experiments to exam the performance of the proposed method under different cloud conditions. In the first set of experiments, we subset a part of the core urban areas for testing, identified a few dates with 0% cloud cover, and randomly selected real-world cloud patterns from a different date as the hypothetical clouds. This set of experiments is to test the proposed method in reconstructing LST under data-rich, clear-sky situations. Test data are from clear-sky LST pixels that are masked out with some real-world cloud pattern from other dates during training. The second set of experiments was conducted over all data, with this validation strategy specifically targeting heavily cloudy situations ($>$80%). On these dates, the valid-observed LST is not clear-sky LST but rather similar to overcast situations because of the high percentage of clouds (the satellite was only able to record data because there were a few gaps between the clouds). Test data are from 20% of the valid-observed LST under the heavily-cloudy situations that were held out during training. This set of experiments is to test the proposed method in partially-cloudy and heavily-cloudy situations. In the third set of experiments, as there is no in-situ LST site within the three cities, we are not able to test the proposed method under completely overcast (100%) or without Landsat overpass situations. Alternatively, we adopted a common approach in the literature [31, 18] and validate the reconstruction results indirectly by comparing its capability to estimate near-surface air temperature data (detailed earlier in section IV-B3). This set of experiments can test LST reconstruction under all situations, but we focus on cloudy situations or without Landsat overpass. These above-mentioned three validation strategies are detailed in the section IV-B.
@@
V-B Results Under Clear-Sky Situations With Real-World Cloud Patterns
(a) NYC
(b) London
(c) Hong Kong
@@ V-B Results Under Clear-Sky Situations With Real-World Cloud Patterns
Figure 7: Scatter plot between reconstructed LST and observed LST under clear-sky situations with real-world cloud patterns. The corresponding reconstruction is shown in Figure 8.
@@
Figure 8: Results under clear-sky situations with real-world cloud patterns. From left to right: LST with real-world cloud patterns, reconstruction, residuals between observed LST and reconstructed LST. From top to bottom: NYC (R2=0.97, RMSE=0.73 K, MAE=0.56 K, Bias=0.01 K); London (R2=0.90, RMSE=0.96 K, MAE=0.74 K, Bias=0.27 K); Hong Kong (R2=0.93, RMSE=0.74 K, MAE=0.56 K, Bias=0.02 K). The corresponding scatter plots are shown in Figure 7.
@@
We first show the results under clear-sky situations. We show the scatter plots between the observed and reconstructed LST in Figure 7, along with the corresponding cloud patterns, reconstructed LST and residuals in Figure 8. The proposed method successfully achieved LST reconstruction reconstructed under clear-sky situations with real-world cloud patterns from other dates. Under 52.0% cloud coverage, the proposed method reached a reconstruction result of R2=0.97, RMSE=0.73 K, MAE=0.56 K, and Bias=-0.01 K for NYC (prediction date: 2023-04-01, cloud pattern date: 2023-06-04). Under 52.03% hypothetical cloud cover, the proposed method reached a reconstruction result of R2=0.90, RMSE=0.96 K, MAE=0.74 K, and Bias=-0.27 K for London (prediction date: 2023-09-06, cloud pattern date: 2023-06-18). Under 56.2% cloud coverage, the proposed method reached a reconstruction result of R2=0.93, RMSE=0.74 K, MAE=0.56 K, and Bias=-0.02 K for Hong Kong (prediction date: 2023-11-17, cloud pattern date: 2023-06-26). All these reconstructions are within the sensor’s accuracy [7]. From the reconstruction images and the residuals (Figure 8), we can see that the reconstruction aligns near perfectly with the observed data.
V-C Results Under Heavily-Cloudy Situations

### v2

@@ IV-A2 London
To demonstrate the generalization of the proposed framework, London is selected as a second study case. Located at 51.5N on the Prime Meridian (0.0E), London is the largest city in the U.K. and Western Europe, with a population of 8.8 million across 1,572 km2 (5,690 residents per km2). London has a temperate oceanic climate, with daily mean temperatures ranging from 5.6 to 19C. Typical nighttime temperatures are around 2.7C, while typical daytime highs are around 23.9C and can be above 28C in summertime. Despite its higher latitude, London experiences a milder annual temperature range compared to NYC, largely due to the influence of the North Atlantic Current. With the exception of a small portion in the southeast, most of London falls within Landsat’s cross-track areas (Path 202 Row 024, Path 201 Row 024). Within the first 16-day circle (Jan 9-24) in 2023, Landsat 8 obtained data on 9th and 18th, Landsat 9 obtained data on 10th and 17th, and then the circle repeats every 16 days, totaling 86 scenes in 2023. The city is divided into fine-scale statistical units (983 Middle Super Output Areas and 4,835 Lower Super Output Areas), most of which are smaller than 1 km2, i.e., finer than the MODIS resolution. These geographical units have mortality and other socioeconomic and health statistics that can be used for social good, but as a result of the lack of high resolution temperature data, existing studies had to settle on using coarse resolution temperature data that limits their capability [14, 59, 60].
IV-A3 Hong Kong
@@
Hong Kong is selected as the third study area to demonstrate the generalization of the proposed framework and method across different continents and climate zones. Hong Kong has a humid subtropical climate, with a daily mean temperature between 16.5–28.9C. Typical winter average daily minimum temperature is 13C in January and average daily maximum is 31.1C in August. Similar to NYC and London, most urban areas are also situated within the Landsat cross-track areas (Path 121 Row 045, Path 122 Row 044). Within the first 16-day circle (Jan 1-16) in 2023, Landsat 9 obtained data on the 1st and 10th, Landsat 8 obtained data on the 2nd and 9th, and then the circle repeats every 16 days, totaling 91 scenes within 2023. As a coastal city in Asia, Hong Kong has a land area of 1,073 km2, housing 7.5 million people (6,989 per km2). The city maintains a large portion (75%) of undeveloped lands, further intensifying population density. As a high-density and subtropical city, Hong Kong faces various extreme climate events, in need of high spatial resolution temperature data to capture local-scale variation due to urban topography, vertical morphology, and proximity to the sea. This is therefore a representative case for assessing the added value of the proposed method in a coastal, high-density, and climate-vulnerable setting.
@@
Figure 5: Meteorological stations of the three cities, overlapped with annual mean temperature and geographical unit boundaries. NYC: 7 stations over 135 merged zip codes; London: 6 stations over 983 Middle Super Output Area; Hong Kong: 23 stations over 452 District Council Ordinary Election Constituency Boundaries.
IV-B Data and Preprocessing
IV-B1 Landsat Land Surface Temperature
@@
We obtained all 2023 Landsat 8 and 9 Collection 2 Level-2 science products and used the analysis-ready LST data from the thermal band (Band 10). NYC is completely located within two Landsat scenes (Path 013 Row 032, Path 014 Row 032), with a total number of 92 scenes. London is completely covered by one scene (Path 201 Row 024), and has its majority covered by another (Path 202 Row 024), with a total number of 86 scenes. Hong Kong has most of its urban areas covered by two scenes (Path 121 Row 045, Path 122 Row 044), with a total number of 91 scenes.
@@
Before conducting the experiments, we went through a series of preprocessing, mainly on cloud masking and removal. We first checked the cloud band from the Level-2 products and masked out all data labeled as unclear. The cloud masks are generated from the Fmask algorithm and are not 100% perfect as noted from data description. We then double-checked visually with the imagery and cloud masks using a $\times3$ split. Within each split, if the pixels were mostly flagged as clear but were in fact covered by clouds, we masked out all the pixels within the split. In some cases, within each split, all pixels within the split appeared clear and most pixels were flagged as clear, but some pixels were flagged as not clear—mostly at coastline—we flagged this split as clear. After preprocessing, we have three types of data: dates under clear-sky situations, dates under partially-cloudy situations, and dates without valid observations (100% cloudy or no Landsat overpass).
IV-B2 ERA5 Reanalysis LST Data
@@
IV-B3 Annual Mean Spectral Reflectance
In order to link valid-observed pixels to cloud-covered pixels, the proposed method DELAG uses Gaussian processes to model the residuals between observed LST and ATC-reconstructed LST. The residuals are related to ground objects, which can be distinguished through spectral reflectance. We obtained the mean surface reflectance from 2022-2024 for four spectral bands (red, green, blue, and near infrared) from the Sentinel-2 products. The data is then resampled and matched with the Landsat ones. We obtained through Sentinel-2 instead of Landsat because Sentinel-2 has a higher revisit frequency, but the influence is negligible as seen from a sensitivity analysis. We also added the x and y coordinates, along with the spectral bands, as the features used in Gaussian processes.
IV-B4 Near-surface Air Temperature Data
There is no in-situ LST site within the three cities, thus a direct comparison with in-situ LST in our study areas is not possible. However, given that most real-world applications require near-surface air temperature instead of LST, and many seamless LST data will be used for generating near-surface air temperature, we can indirectly validate the reconstruction with near-surface air temperature data (details in method Section V-B3). This is an approach commonly used in the existing literature [28, 33]. For this purpose, we obtained air temperature data from meteorological stations. For New York City, we obtained data from 7 stations from the Automated Surface/Weather Observing Systems (ASOS/AWOS) maintained by the National Weather Service. For London, there are 6 meteorological stations. These data were obtained through Meteostat, an open source Python library that archives meteorological data from public sources. For Hong Kong, we obtained air temperature data from 23 stations from the Hong Kong Observatory. The locations of the meteorological stations in the three cities can be found in Fig. 5.
V Methodology
V-A DELAG: Deep Ensemble Learning With Annual Temperature Cycles and Gaussian Processes
@@
Overall, the proposed method DELAG can be expressed as
@@
V-A1 Enhanced Annual Temperature Cycles
@@
V-A2 Gaussian Processes to Model LST Residuals
@@
$\begin{aligned} &\log p\left(\mathcal{T}^{\epsilon}\mid\mathbf{X},\mathbf{K},\Theta\right)\\ &=-\frac{1}{2}\left(\mathcal{T}^{\epsilon}-m\left(\mathbf{X}\right)\right)^{\top}\mathbf{K}^{-1}\left(\mathcal{T}^{\epsilon}-m\left(\mathbf{X}\right)\right)\\ &\;\;\;\;\;\;\;-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log\left(2\pi\right)\;,\end{aligned}$ (6)
@@
V-A3 LST Residuals for Cloud-Covered Pixels
@@
V-A4 Deep Ensemble Learning to Find Model Solutions and Achieve Uncertainty Quantification
Given the complexity of the model as a non-convex function, it is very difficult, if not impossible, to find the global optimized set of parameters. In the PyTorch deep learning framework, we use the Adam optimizer (a variant of stochastic gradient descent) [63] to minimize the L1 loss to find the optimized solutions in the first step. In training the model, the training loss will eventually flatten and fluctuate, but cannot reach a global minimum. In Fig. 6, near the end of training, 334 out of 425 snapshots all fall below the smallest contour level (loss<<0.06). At this phase of the training, all snapshots are acceptable optimized solutions of the model. With each iteration, the parameters update again, resulting another snapshot of model parameters. The disagreement drawn from these snapshots of parameters is large enough, and utilizing the disagreements among these snapshot solutions has become a popular method to achieve ensemble learning or quantify uncertainty in deep learning [64, 65, 66]. Ensemble learning to achieve more robust and reliable prediction is also common in weather predictions and remote sensing [67, 68].
Figure 6: Demo showing the loss surface and model trajectory during training. The x- and y-axes represent unitless, compressed parameter dimensions. As training progresses, the loss decreases substantially, with 334 out of 425 snapshots falling below the lowest contour level (0.06). DELAG leverages these diverse solutions as an ensemble to produce robust predictions with uncertainty quantification.
@@ V-B Validation Strategies
V-B1 Validation With Landsat Clear-Sky Observations Under Real-World Cloud Patterns
@@
V-B2 Validation With Partially-Cloudy Data
@@
V-B3 Validation Via Estimating Near-Surface Air Temperature
@@ V-C Training Setup
The experiments were conducted on Python using the PyTorch framework [72]. For ATC fitting, we used the Adam optimizer with a learning rate of 0.1 for 1200 epochs. The 200 ensemble predictions were obtained through every 4 epochs in the final 800 epochs. For Gaussian processes, we used GPyTorch [62], a popular GP framework built upon PyTorch. We used a mini batch of 1024 and the variational inference with 512 inducing points for fast GP computation. The training was conducted with the Adam optimizer with a learning rate of 0.05 for 50 epochs and then of 0.005 for 10 epochs. The training settings apply to all experiments.
V-D Experimental Setup
@@
We conducted three sets of experiments to examine the performance of the proposed method under different cloud conditions. In the first set of experiments, we subset a part of the core urban areas for testing, identified a few dates with 0% cloud cover, and randomly selected real-world cloud patterns from a different date as the hypothetical clouds. This set of experiments is to test the proposed method in reconstructing LST under data-rich, clear-sky situations. Test data are from clear-sky LST pixels that are masked out with some real-world cloud pattern from other dates during training. The second set of experiments was conducted over all data, with this validation strategy specifically targeting heavily cloudy situations (>>80%). On these dates, the valid-observed LST is not clear-sky LST but rather similar to overcast situations because of the high percentage of clouds (the satellite was only able to record data because there were a few gaps between the clouds). Test data are from 20% of the valid-observed LST under the heavily-cloudy situations that were held out during training. This set of experiments is to test the proposed method in partially-cloudy and heavily-cloudy situations. In the third set of experiments, as there is no in-situ LST site within the three cities, we are not able to test the proposed method under completely overcast (100%) or without Landsat overpass situations. Alternatively, we adopt a common approach in the literature [28, 33] and validate the reconstruction results indirectly by comparing its capability to estimate near-surface air temperature data (detailed earlier in section V-B3). This set of experiments can test LST reconstruction under all situations, but we focus on cloudy situations or without Landsat overpass. These above-mentioned three validation strategies are detailed in the section V-B.
VI Results and Analysis
VI-A Integrating ERA5 Reanalysis Data and Landsat Observations for LST Reconstruction
We first demonstrate that, even within the same ERA5 reanalysis data pixel (11 km), the reconstructed LST can vary substantially due to differences in Landsat observations and the resulting fitted parameters in the eATC models. As shown in Fig. 7, the Landsat pixels, which all correspond to the same ERA5 data pixel, have a wide range of temperature values. The pixel locations are marked in circle in the image, and their relative positions are the same as arranged in the plots. Through different fitted parameters, the eATC models successfully captured the majority of the temporal information. Specifically, the first parameter, representing the annual mean temperature, ranges from 286.91 to 297.54 K. The water pixels (1st row 2nd column, 3rd row 1st column) have lower annual mean temperature, as expected. The water pixels also have a smaller amplitude (2nd parameter), being 2.07 and 2.18 K, compared to land and urban pixels’ 6.72-8.22 K. (The 2nd row 1st column is a land-water boundary pixel with a 3.37 K amplitude.) The linear term on the ERA5 reanalysis data further captures the daily fluctuation through varying the 4th parameter, ranging from 0.99 to 1.03. As the ERA5 data pixel can be considered as the mean of a larger area, its trend is therefore a signal combining all individual Landsat pixels falling within the ERA5 data pixel. Assuming a linear relationship between individual Landsat pixels and the ERA5 data can partially recover Landsat LST data, and, combined with the original three parameters of ATC, from the demo here we can see that the majority of the information is recovered through the eATC model (Fig. 7). The remaining residuals are then recovered through Gaussian processes.
Figure 7: Demo of fitted eATC models for 9 Landsat pixels within a single ERA5 data pixel. The Landsat pixels, marked with circles in the image, are arranged in the subplots to reflect their relative positions.
VI-B Residual Surface Analysis Before Gaussian Processes
If residual surfaces of individual days are related to each other, we can then generalize residual surfaces from days with observations to days without observations. Unfortunately, the residual surface is not identical each day, as the information captured through the eATC model is unbiased and scattered around the observations, meaning that the eATC reconstructed LST can be higher or lower than the actual value. We visualize the residual surfaces from 18 selected days of the NYC data in Fig. 8, showing different patterns across different days. Nonetheless, on the same day, the residual surface is related to ground surface materials (a notable patten is the residuals over water and land areas), which we can use Gaussian processes to model.
Figure 8: Various patterns of the residual surfaces from different days.
@@ VI-B Residual Surface Analysis Before Gaussian Processes
VI-C Results Under Clear-Sky Situations With Real-World Cloud Patterns
@@
Figure 9: Scatter plot of reconstructed versus observed LST under clear-sky conditions, using real-world cloud patterns from other days.
@@
Figure 10: Results under clear-sky conditions using real-world cloud patterns from other days. Columns display: Observed LST (with real-world cloud patterns), Reconstructed LST, and Residuals. Rows represent NYC, London, and Hong Kong.
@@
The results under clear-sky situations with real-world hypothetical cloud patterns from another day are shown in Fig. 9 with scatter plots, and in Fig. 10 with the reconstruction visualizations including the cloud patterns, reconstructed LST, and residuals. The proposed method successfully achieved LST reconstruction under clear-sky situations with real-world cloud patterns from other dates. Under 52.0% hypothetical cloud coverage, the proposed method reached a reconstruction result of R2=0.97, RMSE=0.73 K, MAE=0.56 K, and Bias=-0.01 K for NYC (prediction date: 2023-04-01, cloud pattern date: 2023-06-04). Under 52.03% hypothetical cloud cover, the proposed method reached a reconstruction result of R2=0.90, RMSE=0.96 K, MAE=0.74 K, and Bias=-0.27 K for London (prediction date: 2023-09-06, cloud pattern date: 2023-06-18). Under 56.2% hypothetical cloud coverage, the proposed method reached a reconstruction result of R2=0.93, RMSE=0.74 K, MAE=0.56 K, and Bias=-0.02 K for Hong Kong (prediction date: 2023-11-17, cloud pattern date: 2023-06-26). All these reconstructions are within the sensor’s accuracy [7]. From the reconstruction images and the residuals (Fig. 10), we can see that the reconstruction aligns near perfectly with the observed data.
VI-D Results Under Heavily-Cloudy Situations